# Segmenting Objectiveness and Task-awareness Unknown Region for Autonomous Driving

### v1

@@ Abstract.
copyright: acmcopyrightjournalyear: 2025ccs: Computing methodologies Scene anomaly detection
@@ 1. Introduction
The integration of transformer-based architectures  (Zheng et al., 2021; Cheng et al., 2021, 2022; Ranftl et al., 2021; Strudel et al., 2021)and large-scale pretrained models (Liu et al., 2021; Kirillov et al., 2023; Dosovitskiy et al., 2020) has significantly advanced semantic segmentation in road scene perception. These methods have demonstrated outstanding performance in closed-set environments, characterized by predefined object categories such as vehicles and pedestrians (Cordts et al., 2016). However, real-world environments are inherently open, often containing unexpected out-of-distribution (OOD) objects—ranging from accident debris to rare natural phenomena—that significantly challenge traditional segmentation models. Misclassifying these anomalies as background or assigning them to known categories introduces substantial safety hazards, particularly in autonomous driving applications. For example, failure to detect a fallen tree or a misplaced construction sign can result in severe accidents. Consequently, reliably identifying and segmenting OOD objects without compromising the accuracy of in-distribution (ID) segmentation has emerged as a critical challenge, named as road anomaly detection task, for achieving robust and safe scene understanding (Yang et al., 2021; Bogdoll et al., 2022).
@@
Current road anomaly detection methods have evolved through two main technical routes. The first paradigm primarily relied on reconstruction-based approaches (Lis et al., 2023; Zavrtanik et al., 2020; Xia et al., 2020; Lis et al., 2019; Wang et al., 2018), which utilized generative models to learn the distribution of normal scenes and detected anomalies by identifying regions with high reconstruction errors. These methods demonstrated effectiveness in controlled environments but faced inherent challenges in handling complex real-world scenarios. In recent years, the field has progressively shifted towards score-based methodss (Tian et al., 2022; Di Biase et al., 2021; Liu et al., 2023; Nayal et al., 2023; Rai et al., 2023; Delić et al., 2024), which formulate anomaly detection as a confidence estimation problem. These approaches typically leverage the prediction uncertainty or feature-space distances to compute anomaly scores at the pixel level. Representative works in this direction include RPL (Liu et al., 2023)and RbA (Rai et al., 2023). RPL (Liu et al., 2023) employs residual pattern learning to identify deviations from expected feature distributions, while RbA (Nayal et al., 2023) adopts a multi-head rejection mechanism to treat the result of mask classification as multiple one vs. all classifiers for robust anomaly identification.
@@
Despite significant advancements, previous approaches often treat road anomaly detection as a general anomaly detection problem, where all outliers or out-of-distribution (OOD) objects are treated as anomaly. This leads to two fundamental limitations that hinder their practical deployment in autonomous driving systems, as illustrated in Fig. 1. First, they struggle with incomplete segmentation of partial anomalies. Current methods fail to capture all OOD-specific pixels, particularly when anomalies share partial visual characteristics with in-distribution objects. For example, while the state-of-the-art RbA (Nayal et al., 2023)(Fig. 1, column 1–2) can detect anomalies like damaged tires, it fails to segment them entirely, leaving residual regions misclassified as background. This issue arises because its multi-head rejection mechanism suppresses uncertain pixels but does not explicitly model OOD shape continuity, leading to fragmented masks where parts of the anomaly remain undetected. Second, they suffer from task-agnostic overdetection. Existing methods overlook road scene constraints, leading to false alarms in irrelevant regions. As shown in Fig. 1 (column 3–5), non-critical objects outside drivable areas, such as roadside vegetation and distant buildings, are erroneously flagged as OOD. This occurs because RbA’s(Nayal et al., 2023) threshold-based scoring system prioritizes anomaly likelihOOD without considering autonomous driving task requirement, resulting in overdetection in areas unrelated to navigation safety. These two issues—partial under-segmentation and task-irrelevant overdetection—highlight the need for anomaly detection frameworks that explicitly model OOD objective attributes while incorporating scene constraints to enhance reliability in autonomous driving.
@@
In this paper, we propose Segmenting Objectiveness and Task-awareness (SOTA), a unified framework that integrates semantic feature fusion with scene-understanding guided prompt learning. It consists of two modules: the Semantic Fusion Block (SFB) and the Scene-understanding Guided Prompt-Context Adapter (SG-PCA). For the SFB, the vanilla OOD prediction generated by the pixel-wise segmentor is aligned to the latent vision space through projection and alignment blocks. The fusion of the OOD map with the vision features improves the precision of objectiveness segmentation by refining the detection of OOD-specific attributes and reducing errors associated with partial anomalies. On the other hand, SG-PCA extracts road scene priors (e.g., lane topology, drivable areas) via task-aware aggregation while adapting Erosion and Dilation operations to resolve partial occlusion issues. Subsequently, scene priors and the vanilla OOD prediction are combined to generate task-aware prompts via multi-aware cross-attention. By suppressing navigation-irrelevant anomalies through prompt learning, SG-PCA ensures focus on safety-critical regions, overcoming the environmental constraints inherent in distribution-based methods. Moreover, we introduce parameter-efficient adaptation of SAM’s (Kirillov et al., 2023) mask decoder using Low-Rank Adaptation (LoRA) (Hu et al., 2022), enabling seamless integration of enriched OOD embeddings without additional manual threshold tuning or postprocessing. Our contributions can be summarized as follows:
@@
  • We propose Segmenting Objectiveness and Task-awareness (SOTA), which consists of two key modules—Semantic Fusion Block (SFB) and Scene-understanding Guided Prompt-Context Adapter (SG-PCA)—that together achieve spatial and semantic completeness in anomaly detection.
@@ 2. Related Work
Semantic segmentation is the task of assigning a class label to each pixel in an image, playing a critical role in scene understanding and computer vision (Cordts et al., 2016; Thisanke et al., 2023; Minaee et al., 2020). Early models (Long et al., 2015; Ronneberger et al., 2015; Badrinarayanan et al., 2017; Zhao et al., 2017; Chen et al., 2018) leveraged convolutional neural networks (CNNs) for pixel-wise classification, enabling the extraction of hierarchical features across multiple scales. Building on these foundations, methods such as DeepLab (Chen et al., 2017) introduced dilated convolutions to expand the receptive field without sacrificing resolution, while DeepLabv3+ (Chen et al., 2018) refined segmentation by combining atrous convolution with a decoder module, improving boundary delineation and object accuracy in complex scenes.
More recently, transformer-based architectures (Dosovitskiy et al., 2020; Liu et al., 2021) have demonstrated the ability to model long-range dependencies and global context, offering significant improvements over CNN-based methods. In particular, frameworks like MaskFormer (Cheng et al., 2021) and Mask2222 Former (Cheng et al., 2022), which integrate mask-based attention mechanisms, have unified various segmentation tasks and achieved state-of-the-art performance across multiple benchmarks. Alongside these advancements, prompt-based segmentation approaches—exemplified by the Segment Anything Model (SAM) (Kirillov et al., 2023)—enable more flexible and accurate segmentation by incorporating prompts such as points, boxes, or text. These developments collectively advance the field, driving robust semantic segmentation in increasingly complex real-world scenarios.
However, traditional semantic segmentation models are typically trained on closed domains, meaning they rely on predefined categories of objects and are not designed to handle out-of-distribution (OOD) objects or unknown categories. As a result, they struggle in real-world open-set environments, such as autonomous driving, where new or unexpected objects (e.g., debris, rare natural animals) might appear. For driving safety, semantic segmentation must possess the ability to detect such OOD objects, ensuring reliable scene understanding and robust performance in dynamic, open-world scenarios. This limitation has led to the research focus on road anomaly detection, a critical area aimed at addressing these challenges by enabling detection and segmentation of OOD objects in driving environments.
@@
Anomaly segmentation involves identifying regions in an image that deviate from normal patterns, often for detecting out-of-distribution (OOD) objects or unusual events (Bogdoll et al., 2022; Yang et al., 2021). This task extends traditional semantic segmentation by not only classifying pixels but also recognizing anomalies, making it particularly useful in autonomous driving. Early methods relied on probabilistic models to compute uncertainty scores (e.g., Monte-Carlo dropout, deep ensembles)(Hendrycks and Gimpel, 2016; Lakshminarayanan et al., 2017; Mukhoti and Gal, 2018; Jung et al., 2021) and reconstruction-based approaches(Lis et al., 2023; Zavrtanik et al., 2020; Xia et al., 2020; Lis et al., 2019; Wang et al., 2018), which detect anomalies by measuring the failure in reconstructed input images.
@@
Recent advancements (Delić et al., 2024; Grcic et al., 2023; Nayal et al., 2023; Rai et al., 2023; Liu et al., 2023) have shifted towards more sophisticated models that integrate deep learning with uncertainty estimation. For example, both RbA (Nayal et al., 2023) and Mask2Anomaly (Rai et al., 2023) use mask-class pairs, with RbA deriving anomaly scores for regions not covered by any masks, while Mask2Anomaly transitions from per-pixel to mask classification, employing techniques like global masked attention and mask contrastive learning. Additionally, RPL (Liu et al., 2023) detects OOD pixels while minimizing the impact on inlier segmentation accuracy. Several works also adapt open-set segmentation methods to detect unknown anomalies, leveraging transformer-based architectures or prompt-based models such as the Segment Anything Model (SAM)(Kirillov et al., 2023). These methods, including S2M(Zhao et al., 2024), demonstrate significant performance gains; however, current approaches that use SAM only apply it during inference and do not incorporate training or fine-tuning strategies to optimize SAM for anomaly detection.
@@
However, existing anomaly segmentation methods typically focus primarily on distribution-based considerations, overlooking the unique characteristics of the driving environment. In real-world applications, it is crucial that detected anomalies correspond to meaningful objects or structures, rather than arbitrary regions of the image. Furthermore, the relevance of these anomalies to the driving task, particularly their potential impact on safety, must be considered. In this paper, we propose Segmenting Objectiveness and Task-awareness (SOTA), a method designed to segment anomaly objects with both spatial and semantic awareness. This approach ensures that anomalies are not only identified based on their objectivity but also evaluated in terms of their relevance to the driving task, thereby improving both segmentation accuracy and the system’s ability to prioritize safety-critical regions.
@@
Figure 2. Overview of our SOTA. Our approach comprises two principal modules: (a). The Semantic Fusion Block (SFB), highlighted by the green regions, fuses image features with pixel-level anomaly scores $\tilde{\mathbf{y}}$ to emphasize out-of-distribution cues; (b). The Scene-understanding Guided Prompt Learning (SG-PCA) module is highlighted by the purple regions in our design. First, it applies task-aware aggregation to extract context-specific prompt, denoted as $\mathbf{y^{\prime}}_{t}$. Then, multi-aware cross attention is used to fuse $\mathbf{y^{\prime}}_{t}$ with the raw anomaly scores $\tilde{\mathbf{y}}$, integrating detailed scene context. The fused feature and refined prompt are then input to a LoRA-adapted mask decoder, yielding precise anomaly segmentation $\hat{\mathbf{y}}$.
@@ 3. Methodology
In this work, we address two key limitations of previous anomaly segmentation methods through our Segmenting Objectiveness and Task-awareness (SOTA) framework, which offers a simple yet effective solution.Our approach builds upon a pre-trained pixel-wise OOD detector that generates OOD confidence maps. First, we introduce the Semantic Fusion Block (SFB) , which combines image embeddings from the Segment Anything Model (SAM) (Kirillov et al., 2023) with OOD confidence maps. This fusion ensures that anomalies are effectively highlighted without overshadowing the core in-distribution features. Next, to capture road scene priors, we design the Scene-understanding Guided Prompt-Context Adapter (SG-PCA), which utilizes multi-aware cross-attention to extract and integrate relevant contextual information. Finally, for seamless integration of enriched OOD embeddings without postprocessing, we adopt Low-Rank Adaptation (LoRA) (Hu et al., 2022) in the mask decoder, ensuring a parameter-efficient incorporation of these enriched embeddings.
@@ 3.2. Semantic Fusion Block
Our Semantic Fusion Block is designed to seamlessly combine anomaly cues derived from tokenized OOD features with the primary image embeddings, thereby enhancing the contextual accuracy of anomaly segmentation. Given an outlier-exposed image $\mathbf{x^{oe}}\in\mathbb{R}^{H\times W\times 3}$, the SAM’s image encoder extracts the main image feature map $\ \mathbf{F_{I}}\in\mathbb{R}^{D_{I}\times H^{\prime}\times W^{\prime}},$ where $D_{I}$ is the image feature dimension and $H^{\prime}\times W^{\prime}$ is the spatial resolution. In parallel, a pixel-level segmentor (e.g., RbA) produces an OOD confidence map $\mathbf{\tilde{y}}\in[0,1]^{H\times W}$. To transform this single-channel anomaly confidence map into a rich latent representation, we first project $\tilde{y}$ using a learnable projection function $f_{\mathrm{proj}}(\cdot)$, yielding a token feature $\mathbf{T}\in\mathbb{R}^{D_{T}\times H^{\prime}\times W^{\prime}},$ with $D_{T}$ being the token dimension. To ensure compatibility between the token representation and the image features, we further align $\mathbf{T}$ with $\mathbf{F_{I}}$ via an alignment function $f_{\mathrm{align}}(\cdot)$, resulting in an aligned token
@@ 3.3. Scene-understanding Guided Prompt Context Adapter
To enhance both the precision and the contextual relevance of anomaly segmentation, we propose a process that refines the anomaly prompt by leveraging scene understanding and a context adapter. In this framework, the process is divided into two complementary components: (1) a Task-aware Aggregation module that exploits scene cues (specifically the road region) to generate a semantically meaningful prompt, and (2) a Multi-aware Cross Attention module that employs channel-wise cross attention to dynamically integrate contextual features from the anomaly cues.
@@
In practice, even a well-refined road prompt might not capture the full complexity of out-of-distribution objects, particularly when such objects extend partially into non-drivable regions due to perspective or occlusion. To tackle this limitation, our second stage introduces a channel-wise cross attention that employs channel-wise cross attention to more fully integrate the raw anomaly signals with the scene-informed prompt.
Specifically, we begin by projecting both the task-guided anomaly map $\tilde{\mathbf{y}}_{t}$ and the original anomaly confidence map $\tilde{\mathbf{y}}$ into a shared embedding space. This projection is achieved via learnable 1×1 convolutional layers—denoted as $f_{Q}$, $f_{K}$, and $f_{V}$—which produce the query, key, and value representations, respectively:
(8) $\mathbf{Q}$ $=f_{Q}(\tilde{\mathbf{y}}_{t}),\quad\mathbf{K}=f_{K}(\tilde{\mathbf{y}}),\quad\mathbf{V}=f_{V}(\tilde{\mathbf{y}}),$
with $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{d\times H^{\prime}\times W^{\prime}}$ and $d$ being the common embedding dimension (and $H^{\prime}\times W^{\prime}$ the spatial resolution after appropriate resizing).We treat every single channel attention map as a token.
The channel-wise cross attention mechanism then computes an attention matrix $\mathbf{A}\in\mathbb{R}^{d\times d}$ by aligning the task-guided features with the raw anomaly features:
@@
with the softmax function applied over the channel dimension. This attention matrix $\mathbf{A}$ captures the inter-channel correlations, reflecting how different channels of the refined prompt relate to the original anomaly map.
@@
The resulting embedding fuses the spatial constraints from the refined road prompt with comprehensive anomaly information, thereby enabling the mask decoder to achieve robust and precise segmentation of out-of-distribution objects. This anomaly detection in scene context through road extraction and enhancing it with channel-wise cross attention—ensures robustness to variations such as perspective-induced partial object appearances while maintaining high segmentation accuracy.

### v2

@@ Abstract.
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Ireland.booktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelandisbn: 979-8-4007-2035-2/2025/10doi: XXXXXXX.XXXXXXXccs: Computing methodologies Scene anomaly detection
@@ 1. Introduction
The integration of transformer-based architectures  (Zheng et al., 2021; Cheng et al., 2021, 2022; Ranftl et al., 2021; Strudel et al., 2021)and large-scale pretrained models (Liu et al., 2021; Kirillov et al., 2023; Dosovitskiy et al., 2020) has significantly advanced semantic segmentation in road scene perception. These methods have shown outstanding performance in closed-set environments, characterized by predefined object categories like vehicles and pedestrians (Cordts et al., 2016). However, real-world environments are inherently open, often containing unexpected out-of-distribution (OOD) objects—from accident debris to rare natural phenomena—that significantly challenge traditional segmentation models. Misclassifying these anomalies as background or assigning them to known categories introduces substantial safety hazards, particularly in autonomous driving. For instance, failing to detect a fallen tree or a misplaced construction sign can result in severe accidents. Consequently, reliably identifying and segmenting OOD objects without compromising in-distribution (ID) segmentation accuracy has emerged as a critical challenge, termed road anomaly detection, for robust and safe scene understanding (Yang et al., 2021; Bogdoll et al., 2022).
@@
Current road anomaly detection methods follow two main technical routes. The first paradigm relies primarily on reconstruction-based approaches (Lis et al., 2023; Zavrtanik et al., 2020; Xia et al., 2020; Lis et al., 2019; Wang et al., 2018), which use generative models to learn normal scene distributions and detect anomalies via high reconstruction errors. These methods proved effective in controlled environments but faced inherent challenges in complex real-world scenarios. Recently, the field has shifted towards score-based methods (Tian et al., 2022; Di Biase et al., 2021; Liu et al., 2023; Nayal et al., 2023; Rai et al., 2023; Delić et al., 2024), framing anomaly detection as a confidence estimation problem. These approaches typically leverage prediction uncertainty or feature-space distances to compute pixel-level anomaly scores. Representative works include RPL (Liu et al., 2023)and RbA (Rai et al., 2023). RPL (Liu et al., 2023) uses residual pattern learning to identify deviations from expected feature distributions, while RbA (Nayal et al., 2023) employs a multi-head rejection mechanism, treating mask classification results as multiple one-vs-all classifiers for robust anomaly identification.
@@
Despite significant advancements, prior approaches often treat road anomaly detection as a general anomaly detection problem, where all outliers are deemed anomalous. This leads to two fundamental limitations hindering their practical deployment in autonomous driving systems, as shown in Fig. 1. First, they struggle with incomplete segmentation of partial anomalies. Current methods fail to capture all OOD-specific pixels, especially when anomalies share partial visual characteristics with in-distribution objects. For example, state-of-the-art RbA (Nayal et al., 2023) (Fig. 1, column 1–2) detects anomalies like damaged tires but fails to segment them entirely, leaving residual regions misclassified as background. This issue arises because its multi-head rejection mechanism suppresses uncertain pixels without explicitly modeling OOD shape continuity, leading to fragmented masks with undetected anomaly parts. Second, they suffer from task-agnostic overdetection. Existing methods overlook road scene constraints, leading to false alarms in irrelevant regions. As shown in Fig. 1 (column 3–5), non-critical objects outside drivable areas, such as roadside vegetation and distant buildings, are erroneously flagged as OOD. This happens because RbA’s(Nayal et al., 2023) threshold-based scoring system prioritizes anomaly likelihood without considering autonomous driving task requirements, resulting in overdetection in areas unrelated to navigation safety. These two issues—partial under-segmentation and task-irrelevant overdetection—highlight the need for anomaly detection frameworks that explicitly model OOD objective attributes while incorporating scene constraints to enhance autonomous driving reliability.
@@
In this paper, we propose Segmenting Objectiveness and Task-awareness (SOTA), a unified framework integrating semantic feature fusion with scene-understanding guided prompt learning. It comprises two modules: the Semantic Fusion Block (SFB) and the Scene-understanding Guided Prompt-Context Adapter (SG-PCA). For SFB, the pixel-wise segmentor’s vanilla OOD prediction is aligned to the latent vision space via projection and alignment blocks. Fusing the OOD map with vision features improves objectiveness segmentation precision by refining OOD-specific attribute detection and reducing partial anomaly errors. SG-PCA, conversely, extracts road scene priors (e.g., lane topology, drivable areas) through task-aware aggregation, adapting Erosion and Dilation operations to resolve partial occlusion issues. Subsequently, scene priors and the vanilla OOD prediction combine to generate task-aware prompts via multi-aware cross-attention. By suppressing navigation-irrelevant anomalies through prompt learning, SG-PCA ensures focus on safety-critical regions, overcoming environmental constraints inherent in distribution-based methods. Moreover, we introduce parameter-efficient adaptation of SAM’s (Kirillov et al., 2023) mask decoder using Low-Rank Adaptation (LoRA) (Hu et al., 2022), enabling seamless integration of enriched OOD embeddings without additional manual threshold tuning or postprocessing. Our contributions are summarized as follows:
@@
  • We propose Segmenting Objectiveness and Task-awareness (SOTA), comprising two key modules—Semantic Fusion Block (SFB) and Scene-understanding Guided Prompt-Context Adapter (SG-PCA)—that together achieve spatial and semantic completeness in anomaly detection.
@@ 2. Related Work
Semantic segmentation, the task of assigning a class label to each pixel in an image, is critical in scene understanding and computer vision (Cordts et al., 2016; Thisanke et al., 2023; Minaee et al., 2020). Early models (Long et al., 2015; Ronneberger et al., 2015; Badrinarayanan et al., 2017; Zhao et al., 2017; Chen et al., 2018) leveraged convolutional neural networks (CNNs) for pixel-wise classification, extracting hierarchical features across scales. Building on this, methods like DeepLab (Chen et al., 2017) used dilated convolutions for expanded receptive fields without sacrificing resolution, while DeepLabv3+ (Chen et al., 2018) refined segmentation by combining atrous convolution with a decoder, improving boundary delineation and object accuracy.
More recently, transformer-based architectures (Dosovitskiy et al., 2020; Liu et al., 2021) have shown the ability to model long-range dependencies and global context, significantly improving over CNNs. Frameworks like MaskFormer (Cheng et al., 2021) and Mask2Former (Cheng et al., 2022), integrating mask-based attention, have unified various segmentation tasks and achieved state-of-the-art performance. Alongside these, prompt-based segmentation, exemplified by the Segment Anything Model (SAM) (Kirillov et al., 2023), enables more flexible and accurate segmentation via prompts like points, boxes, or text. These developments collectively advance the field, driving robust semantic segmentation in increasingly complex real-world scenarios.
However, traditional semantic segmentation models, trained on closed domains, rely on predefined categories. They are not designed for out-of-distribution (OOD) objects or unknown categories. Thus, they struggle in open-set environments like autonomous driving, where new or unexpected objects (e.g., debris, rare animals) may appear. For driving safety, semantic segmentation must detect such OOD objects, ensuring reliable scene understanding and robust performance in dynamic, open-world scenarios. This limitation has shifted research focus to road anomaly detection, a critical area addressing these challenges by enabling OOD object detection and segmentation in driving environments.
@@
Anomaly segmentation involves identifying image regions deviating from normal patterns, often for detecting out-of-distribution (OOD) objects or unusual events (Bogdoll et al., 2022; Yang et al., 2021). This task extends traditional semantic segmentation by classifying pixels and recognizing anomalies, making it particularly useful in autonomous driving. Early methods relied on probabilistic models to compute uncertainty scores (e.g., Monte-Carlo dropout, deep ensembles)(Hendrycks and Gimpel, 2016; Lakshminarayanan et al., 2017; Mukhoti and Gal, 2018; Jung et al., 2021) and reconstruction-based approaches(Lis et al., 2023; Zavrtanik et al., 2020; Xia et al., 2020; Lis et al., 2019; Wang et al., 2018), which detect anomalies by measuring reconstruction failure.
@@
Recent advancements (Delić et al., 2024; Grcic et al., 2023; Nayal et al., 2023; Rai et al., 2023; Liu et al., 2023, 2025) have shifted towards more sophisticated models integrating deep learning with uncertainty estimation. For example, both RbA (Nayal et al., 2023) and Mask2Anomaly (Rai et al., 2023) use mask-class pairs; RbA derives anomaly scores for regions not covered by masks, while Mask2Anomaly transitions from per-pixel to mask classification, employing global masked attention and mask contrastive learning. Additionally, RPL (Liu et al., 2023) detects OOD pixels while minimizing impact on inlier segmentation accuracy. Several works also adapt open-set segmentation methods to detect unknown anomalies, leveraging transformer-based architectures or prompt-based models like SAM (Kirillov et al., 2023). These methods, including S2M (Zhao et al., 2024), show significant performance gains; however, current approaches using SAM only apply it during inference and lack training or fine-tuning strategies to optimize SAM for anomaly detection.
@@
However, existing anomaly segmentation methods typically focus on distribution-based considerations, overlooking the unique characteristics of the driving environment. In real-world applications, detected anomalies must correspond to meaningful objects or structures, not arbitrary regions. Furthermore, their relevance to the driving task, particularly safety impact, must be considered. In this paper, we propose Segmenting Objectiveness and Task-awareness (SOTA), a method designed to segment anomaly objects with both spatial and semantic awareness. This approach ensures anomalies are identified based on objectivity and evaluated for driving task relevance, thereby improving segmentation accuracy and the system’s ability to prioritize safety-critical regions.
@@
Figure 2. Overview of our SOTA. Our approach comprises two principal modules: (a). The Semantic Fusion Block (SFB), highlighted by the green regions, fuses image features with pixel-level anomaly scores $\tilde{\mathbf{y}}$ to emphasize out-of-distribution cues; (b). The Scene-understanding Guided Prompt Learning (SG-PCA) module is highlighted by the purple regions in our design. First, it applies task-aware aggregation to extract context-specific prompt, denoted as $\mathbf{y^{\prime}}_{t}$. Then, hadamard product and multi-aware cross attention is used to fuse $\mathbf{y^{\prime}}_{t}$ with the raw anomaly scores $\tilde{\mathbf{y}}$, integrating detailed scene context. The fused feature and refined prompt are then input to a LoRA-adapted mask decoder, yielding precise anomaly segmentation $\hat{\mathbf{y}}$.
@@ 3. Methodology
Our approach builds upon a pre-trained pixel-wise OOD detector that generates OOD confidence maps. First, we introduce the Semantic Fusion Block (SFB) , which combines image embeddings from the Segment Anything Model (SAM) (Kirillov et al., 2023) with OOD confidence maps. This fusion ensures that anomalies are effectively highlighted without overshadowing the core in-distribution features. Next, to capture road scene priors, we design the Scene-understanding Guided Prompt-Context Adapter (SG-PCA), which utilizes multi-aware cross-attention to extract and integrate relevant contextual information. Finally, for seamless integration of enriched OOD embeddings without postprocessing, we adopt Low-Rank Adaptation (LoRA) (Hu et al., 2022) in the mask decoder, ensuring a parameter-efficient incorporation of these enriched embeddings.
@@ 3.2. Semantic Fusion Block
[…]
@@ 3.3. Scene-understanding Guided Prompt Context Adapter
To enhance both the precision and the contextual relevance of anomaly segmentation, we propose a process that refines the anomaly prompt by leveraging scene understanding and a context adapter. In this framework, the process is divided into two complementary components: (1) a Task-aware Aggregation module that exploits scene cues (specifically the road region) to generate a semantically meaningful prompt, and (2) a Multi-aware Cross Attention module that employs road-wise cross attention to dynamically integrate contextual features from the anomaly cues.
@@
In practice, even a well-refined road prompt might not capture the full complexity of out-of-distribution objects, particularly when such objects extend partially into non-drivable regions due to perspective or occlusion. To tackle this limitation, our second stage introduces a road-wise cross attention that employs cross attention to more fully integrate the raw anomaly signals with the scene-informed prompt.
Specifically, we begin by projecting both the task-guided anomaly map $\tilde{\mathbf{y}}_{t}$ and the original anomaly confidence map $\tilde{\mathbf{y}}$ into a shared embedding space. This projection is achieved via learnable 1×1 convolutional layers denoted as $f_{Q}$, $f_{K}$, and $f_{V}$, which produce the query, key, and value representations, respectively:
[…]
with $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{d\times H^{\prime}\times W^{\prime}}$ and $d$ being the common embedding dimension (and $H^{\prime}\times W^{\prime}$ the spatial resolution after appropriate resizing).
The road-wise cross attention mechanism then computes an attention matrix $\mathbf{A}\in\mathbb{R}^{d\times d}$ by aligning the task-guided features with the raw anomaly features:
@@
with the softmax function applied over the channel dimension. This attention matrix $\mathbf{A}$ captures the inter correlations, reflecting how road area of the refined prompt relate to the original anomaly map.
@@
The resulting embedding fuses the spatial constraints from the refined road prompt with comprehensive anomaly information, thereby enabling the mask decoder to achieve robust and precise segmentation of out-of-distribution objects. This anomaly detection in scene context through road extraction and enhancing it with road-wise cross attention ensures robustness to variations such as perspective-induced partial object appearances while maintaining high segmentation accuracy.